# DualPM: Dual Posed-Canonical Point Maps for 3D Shape and Pose Reconstruction

### v4

@@ Abstract
Figure 1: Left: We map an image of an object to its Dual Point Maps (DualPMs), a pair of point maps $P$, defined in a camera space, and $Q$, defined in a canonical space where the object has a neutral pose. The pose is thus given by the flow $P-Q$. Right: The DualPMs are easy to predict with a neural network, enabling effective 3D object reconstruction and facilitating geometric tasks like detecting 3D keypoints and fitting a 3D skeleton. For visualization, we color each point with its coordinate in the canonical point maps.
@@ 2 Related work
Figure 3: Method overview. We preprocess the input image $I$ by obtaining a segmentation mask $M$ and image features $F$ using pretrained networks. Subsequently, we predict the layered canonical point map $\{Q_{i}^{*}\}_{i=1}^{2K}$. Conditioned on $\{Q_{i}^{*}\}_{i=1}^{2K}$, we predict the layered posed point map $\{P_{i}^{*}\}_{i=1}^{2K}$, along with the associated opacity maps $\{\sigma_{i}^{*}\}_{i=3}^{2K}$. Both layered point maps are then decoded to produce the canonical point map $Q$ and the posed point map $P$. The training process is supervised using the ground truth point maps and opacity maps.
@@ 4 Experiments
Method PCK (%) Chamfer Distance (cm)
Horse Cow Sheep Real-Sized Normalized
Horse Cow Sheep Horse Cow Sheep
A-CSM [26] 32.9 26.3 28.6 7.60 ± $\pm\ 3.07$ 6.71 ± $\pm\ 1.81$ 2.84 ± $\pm\ 0.77$ 2.73 ± $\pm\ 1.13$ 2.35 ± $\pm\ 0.68$ 2.48 ± $\pm\ 0.70$
MagicPony [72] 42.9 42.5 41.2 7.19 ± $\pm\ 2.35$ 7.22 ± $\pm\ 1.53$ 3.43 ± $\pm\ 0.73$ 2.58 ± $\pm\ 0.80$ 2.53 ± $\pm\ 0.59$ 3.00 ± $\pm\ 0.68$
Farm3D [18] 49.1 40.2 36.1 7.65 ± $\pm\ 2.21$ 6.91 ± $\pm\ 1.49$ 3.79 ± $\pm\ 0.55$ 2.76 ± $\pm\ 0.83$ 2.41 ± $\pm\ 0.54$ 3.31 ± $\pm\ 0.49$
3D-Fauna [33] 53.9 8.69 ± $\pm\ 2.38$ 9.19 ± $\pm\ 2.40$ 3.51 ± $\pm\ 0.88$ 3.13 ± $\pm\ 0.85$ 3.20 ± $\pm\ 0.80$ 3.06 ± $\pm\ 0.76$
Ours 73.2 66.86 66.82 3.13 ± $\pm\ 2.13$ 2.63 ± $\pm\ 1.23$ 1.52 ± $\pm\ 0.88$ 1.11 ± $\pm\ 0.73$ 0.92 ± $\pm\ 0.43$ 1.33 ± $\pm\ 0.78$
Table 1: Quantitative evaluation. We evaluate on PASCAL VOC, reporting PCK@0.1 (higher is better $\uparrow$), and on Animodel [18], reporting the bi-directional Chamfer Distance in centimeters (lower is better $\downarrow$). Our model, trained solely on data from one or two models per category, outperforms other state-of-the-art approaches.
@@ 4.2 Quantitative evaluation
We also measure the bi-directional Chamfer distance between the predicted point maps and the ground truth 3D shape on the Animodel dataset introduced in Farm3D [18]. This dataset comprises renderings of 3D models of horses, cows, and sheep in various poses. To ensure a fair comparison, we train a separate model on data that excludes the poses used in this benchmark. The performance of our method and others is reported in Tab. 1.
@@ 4.4 Ablation study
Method Chamfer Distance (cm) $\downarrow$
Real-Sized Normalized
@@
(supervised 3D-Fauna [33]) 4.96 ± $\pm\ 1.66$ 1.78 ± $\pm\ 0.59$
Ours - 8-layers 3.13 ± $\pm\ 1.98$ 1.11 ± $\pm\ 0.69$
Ours - 4-layers 3.13 ± $\pm\ 2.13$ 1.11 ± $\pm\ 0.73$
   $P$ conditioned on $F$ and $Q$ 3.29 ± $\pm\ 2.26$ 1.17 ± $\pm\ 0.80$
   $P$ conditioned on $F$ 3.70 ± $\pm\ 1.79$ 1.32 ± $\pm\ 0.61$
Table 2: Ablation study on Animodel Horses. We report the bi-directional Chamfer Distance in centimeters (lower is better $\downarrow$) for Animodel Horses. The evaluation follows the protocol as Tab. 1.
@@ Appendix C Technical details
We obtain the segmentation mask $M$ using the Segment Anything method [23]. The feature extractor $\Psi$ is based on [86] which combines pre-trained DINOv2 [45] and StableDiffusion [49] networks. Training image features are reduced to a 64-dimensional space using PCA following [70]. The dual point map predictors $\Phi_{Q}$ and $\Phi_{P}$ leverage a convolutional U-Net architecture based on [55], comprising two blocks each and trained from scratch. We predict $N=4$ layers for layered amodal point maps as more have little effect on the performance (Sec. 4.4), likely due to the low frequency of multiple self-occlusions in our datasets. The number of layers can be easily increased should the data require it. The output resolution of the layered point maps is set to 160$\times$160.

### v5

@@ Abstract
Figure 1: Left: We map an image of an object to its Dual Point Maps (DualPMs), a pair of point maps $P$, defined in a camera space, and $Q$, defined in a canonical space where the object has a neutral pose. The pose is thus given by the flow PQP-Q. Right: The DualPMs are easy to predict with a neural network, enabling effective 3D object reconstruction and facilitating geometric tasks like detecting 3D keypoints and fitting a 3D skeleton. For visualization, we color each point with its coordinate in the canonical point maps.
@@ 2 Related work
Figure 3: Method overview. We preprocess the input image II by obtaining a segmentation mask $M$ and image features $F$ using pretrained networks. Subsequently, we predict the layered canonical point map $\{Q_{i}^{*}\}_{i=1}^{2K}$. Conditioned on $\{Q_{i}^{*}\}_{i=1}^{2K}$, we predict the layered posed point map $\{P_{i}^{*}\}_{i=1}^{2K}$, along with the associated opacity maps $\{\sigma_{i}^{*}\}_{i=3}^{2K}$. Both layered point maps are then decoded to produce the canonical point map $Q$ and the posed point map $P$. The training process is supervised using the ground truth point maps and opacity maps.
@@ 4 Experiments
Method PCK (%) RMS Chamfer Distance (cm) Model-view RMS Chamfer Distance (cm)
Horse Cow Sheep Horse Cow Sheep Horse Cow Sheep
A-CSM [26] 32.9 26.3 28.6 11.75 ± $\pm\ 3.83$ 9.52 ± $\pm\ 2.41$ 9.24 ± $\pm\ 2.40$ 38.13 ± $\pm\ 13.89$ 33.51 ± $\pm\ 11.52$ 29.04 ± $\pm\ 9.35$
MagicPony [72] 42.9 42.5 41.2 11.19 ± $\pm\ 3.08$ 10.29 ± $\pm\ 2.08$ 20.82 ± $\pm\ 13.04$ 25.39 ± $\pm\ 13.43$
Farm3D [18] 49.1 40.2 36.1 11.34 ± $\pm\ 3.22$ 9.63 ± $\pm\ 2.02$ 11.01 ± $\pm\ 1.87$ 29.52 ± $\pm\ 15.73$ 21.34 ± $\pm\ 12.85$ 21.52 ± $\pm\ 9.84$
3D-Fauna [33] 53.9 11.86 ± $\pm\ 3.03$ 10.54 ± $\pm\ 2.26$ 9.61 ± $\pm\ 2.15$ 15.70 ± $\pm\ 6.82$ 14.08 ± $\pm\ 4.20$ 12.24 ± $\pm\ 3.17$
Trellis [73] 6.93 ± $\pm\ 4.13$ 6.80 ± $\pm\ 3.24$ 5.91 ± $\pm\ 2.93$ 36.82 ± $\pm\ 16.02$ 26.54 ± $\pm\ 13.14$ 26.56 ± $\pm\ 12.06$
DualPM (Ours) 73.2 66.86 66.82 4.30 ± $\pm\ 1.50$ 3.18 ± $\pm\ 1.06$ 3.30 ± $\pm\ 1.20$ 5.49 ± $\pm\ 1.75$ 4.03 ± $\pm\ 2.11$ 4.22 ± $\pm\ 2.18$
Table 1: Quantitative evaluation. We evaluate on PASCAL VOC, reporting PCK@0.1 (higher is better $\uparrow$), and on Animodel-Points, a derivative of Animodel  [18] for evaluating point clouds, reporting the Root-mean-square-error of bi-directional Chamfer Distance in centimeters (lower is better $\downarrow$). Point clouds are registered using the iterative closest points algorithm with bidirectional chamfer distance as its cost function. We report both the results with and without rotation alignment. Our model, trained solely on data from one or two models per category, outperforms other state-of-the-art approaches, and has superior camera alignment.
@@ 4.2 Quantitative evaluation
We also measure the bi-directional Chamfer distance between the predicted point maps and the ground truth 3D shape on the Animodel dataset introduced in Farm3D [18]. This dataset comprises renderings of 3D models of horses, cows, and sheep in various poses. To ensure a fair comparison, we train a separate model on data that excludes the poses used in this benchmark. The performance of our method and others is reported in Tab. 1111Amendment: the Animodel evaluation was not suitable for evaluating point clouds, and we have updated our table with an improved benchmark. The original published table is available in the appendix in Tab. A4. Benchmark details are available in App. D.
@@ 4.4 Ablation study
Method RMS Chamfer Distance (cm) $\downarrow$
@@
(supervised 3D-Fauna [33]) 6.79 $\pm$ 2.47
Our method
   8-layers 4.45 $\pm$ 1.61
   4-layers 4.30 $\pm$ 1.50
Ablations on 4-layer model
   $P$ conditioned on $F$ and $Q$ 4.54 $\pm$ 1.60
   $P$ conditioned on $F$ 4.48 $\pm$ 1.46
Table 2: Ablation study on Animodel-points horses category. We report the root-mean-square bi-directional Chamfer Distance in centimetres (lower is better $\downarrow$) for the horse category of Animodel-points. The evaluation follows the protocol as Tab. 1.
@@ Appendix C Technical details
We obtain the segmentation mask $M$ using the Segment Anything method [23]. The feature extractor $\Psi$ is based on [86] which combines pre-trained DINOv2 [45] and StableDiffusion [49] networks. Training image features are reduced to a 64-dimensional space using PCA following [70]. The dual point map predictors $\Phi_{Q}$ and $\Phi_{P}$ leverage a convolutional U-Net architecture based on [55], comprising two blocks each and trained from scratch. We predict $N=4$ layers for layered amodal point maps as more have little effect on the performance (Sec. 4.4), likely due to the low frequency of multiple self-occlusions in our datasets. The number of layers can be easily increased should the data require it. The output resolution of the layered point maps is set to $\times160$.
@@
Appendix D Benchmark details
We introduce Animodel-Points, a benchmark specifically designed to address the limitations of using the original Animodel benchmark for evaluating point cloud generation models. Previous evaluations on Animodel were suboptimal for two primary reasons: the benchmark was designed for meshes, not point clouds, and its evaluation backend (trimesh.registration) [^2] Trimesh: A Python library for loading and using triangular meshes. https://github.com/mikedh/trimesh only optimizes a one-sided Chamfer distance.
Animodel-Points resolves these issues by establishing a new protocol that directly evaluates point clouds using a more robust metric. The key improvements are:
  • Model-view coordinates: Targets are defined in model-view coordinates up to a depth offset, which allows for the additional evaluation of camera alignment performance.
  • Bidirectional Chamfer distance: The evaluation protocol minimizes the mean squared bidirectional Chamfer distance, aligning the optimization objective with the reported cost for a fairer assessment.
  • Standardized point cloud processing: A clear pre-processing and evaluation pipeline ensures consistency and comparability of results.
Methodology
The methodology is divided into two stages: preparing the ground truth data and the protocol for evaluating a prediction.
Data pre-processing
To convert the original Animodel meshes into the Animodel-Points format, we apply the following steps:
  1. 1.
    Coordinate transformation: The source mesh is transformed into model-view coordinates.
  2. 2.
    Uniform scaling: The mesh is uniformly scaled by a factor of $V^{-1/3}$, where $V$ is the mesh volume, to normalize its size.
  3. 3.
    Point sampling: 20,000 points are uniformly sampled from the surface of the scaled mesh to serve as the ground truth target.
Evaluation protocol
For a given generated point cloud, the following evaluation steps are performed:
  1. 1.
    Resampling: If the input point cloud does not contain 20,000 points, it is resampled to this size.
  2. 2.
    Rotational ambiguity handling: The 20,000 point sample is duplicated, and the duplicate is rotated 180 degrees around its vertical axis.
  3. 3.
    Subsampling for fitting: A subset of 10,000 points is sampled from the input to be used in the alignment process.
  4. 4.
    Alignment: Both the original and rotated samples are aligned to the ground truth target using the Iterative Closest Point (ICP) algorithm. The process runs for a maximum of 200 steps or until convergence, using the MSE bidirectional Chamfer distance as the objective function.
  5. 5.
    Final score: The lower of the two costs from the alignment steps is reported as the final score.
Chamfer distance performance is evaluated by providing scale, rotation, and translation degrees of freedom to the transformation estimation. Model-view chamfer distance is evaluated by restricting the degrees of freedom to just scale and translation.
Published results
Method Chamfer Distance (cm)
Real-Sized Normalized
Horse Cow Sheep Horse Cow Sheep
A-CSM [26] 7.60 ± $\pm\ 3.07$ 6.71 ± $\pm\ 1.81$ 2.84 ± $\pm\ 0.77$ 2.73 ± $\pm\ 1.13$ 2.35 ± $\pm\ 0.68$ 2.48 ± $\pm\ 0.70$
MagicPony [72] 7.19 ± $\pm\ 2.35$ 7.22 ± $\pm\ 1.53$ 3.43 ± $\pm\ 0.73$ 2.58 ± $\pm\ 0.80$ 2.53 ± $\pm\ 0.59$ 3.00 ± $\pm\ 0.68$
Farm3D [18] 7.65 ± $\pm\ 2.21$ 6.91 ± $\pm\ 1.49$ 3.79 ± $\pm\ 0.55$ 2.76 ± $\pm\ 0.83$ 2.41 ± $\pm\ 0.54$ 3.31 ± $\pm\ 0.49$
3D-Fauna [33] 8.69 ± $\pm\ 2.38$ 9.19 ± $\pm\ 2.40$ 3.51 ± $\pm\ 0.88$ 3.13 ± $\pm\ 0.85$ 3.20 ± $\pm\ 0.80$ 3.06 ± $\pm\ 0.76$
Ours 3.13 ± $\pm\ 2.13$ 2.63 ± $\pm\ 1.23$ 1.52 ± $\pm\ 0.88$ 1.11 ± $\pm\ 0.73$ 0.92 ± $\pm\ 0.43$ 1.33 ± $\pm\ 0.78$
Table A4: Evaluation on Animodel. We evaluate on Animodel [18], reporting the bi-directional Chamfer Distance in centimeters (lower is better $\downarrow$). Ours include results from models trained on data from each category.
Previously published results were reported on the Animodel benchmark and are available in this table Tab. A4.